The frequency of a photon associated with an energy of 3.31ev is 

Grade:

1 Answers

Arun
5 years ago
Dear student
 
 as we knowE = h\nu3.31 * 1.6 * 10 –19 = 6.626 * 10 –34 * \nuhence\nu = 0.8 * 1015 Hz